OpenStudy (anonymous):

if n!/(n-1)!=5 ,, then (n-2)! =? plz do help me ... ... i dont know how to do this question ...

OpenStudy (amistre64):

what is n! defined as? what is (n-1)! defined as?

OpenStudy (amistre64):

'expand' the factorials and youll be able to solve for n

OpenStudy (cj49):

if n=5 then, 5!/(5-1)!=5 therefore (5-2)!=3! i.e 3x2x1=6

OpenStudy (anonymous):

@cj49 from where 3(2)(1) came ?

OpenStudy (cj49):

the factorial of a non-negative integer n, denoted by n!, is the product of all positive integers less than or equal to n.

OpenStudy (cj49):

for eg 5!=5x4x3x2x1=120

OpenStudy (anonymous):

ahh :O

OpenStudy (anonymous):

can u tell me what is this ! ??

OpenStudy (cj49):

!=factorial

OpenStudy (cj49):

http://www.youtube.com/watch?v=4j66DS_XTSo

hartnn (hartnn):

\(\Large n! = n\times (n-1)\times (n-2)\times ... \times 3\times 2\times 1\) example \(\Large 7! = 7\times6\times5\times4\times3\times2\times1\)

hartnn (hartnn):

\(\Large n! = n\times [(n-1)\times (n-2)\times ... \times 3\times 2\times 1] \\ \Large = n \times (n-1)!\)

OpenStudy (anonymous):

thank u so much @cj49 & @hartnn thsnkx alot for helping me :D

OpenStudy (cj49):

np c:

OpenStudy (anonymous):

:) `